a.Explain Keynesian theory of consumption with its important features.
b.Suppose,the saving of an economy is given by S=-10+0.2Y and autonomous investment I=$500million.
Find:
1.the equilibrium level of income;
2.the level of consumption;
3.the new levels of income and consumption if investment increases permanently by $500 million.
1
Expert's answer
2015-01-13T11:24:28-0500
A. Keynesian theory of consumption is called the theory of absolute income. It is based on the following assumptions:
1. The level of consumption depends only on the magnitude of current disposable income: C = C (Yd), and this relationship is positive, i.e., with an increase in disposable income consumption increases, but
2. In the economy a psychological law acts according to which "people tend to increase their consumption with income growth, but to a lesser extent than income increases." This is due to the fact that because disposable income is divided into consumption and savings:
Yd = C + S,
3. then the growth of disposable income both consumption and savings increase.
B. The main formula is: Y = C + I. Knowing the formula of consumption, it is possible to derive a formula on savings: S = Y- C.
Thus, S = I. The equilibrium level of income is: -10+0.2Y = 500 000. Y = $2.5 billion
The level of consumption is: C = 2500 - 500 = $2 billion.
The new level of income due to investment increase is: Y = $5 billion.
The new level of consumption due to investment increase is: C = $4 billion.